### v3

# M. C. Rodriguez

Abstract
We will study in details the full scalar potential of the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model when all the scalars fields belong to sextet and anti-sextet representation obtain VEV. We will present numerical predictions for neutral scalars of this model, we will show that the mass of the light scalar of CP par is $125.5$ GeV while the light pseudoscalar have mass bigger than $150$ GeV.
Spontaneous breaking of a global symmetry in the Minimum Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model.
M. C. Rodriguez
Grupo de Física Teórica e Matemática Física
Departamento de Física
Universidade Federal Rural do Rio de Janeiro - UFRRJ
BR 465 Km 7, 23890-000 Seropédica - RJ
Brasil
PACS number(s): 12.60. Jv
Keywords: Supersymmetric models
1 Introduction
In the context of the Standard Model (SM) [1, 2], we introduce only left-handed neutrinos in agreement with the two component theory of neutrino of Landau [3], Lee and Yang [4] and Salam [5] from 1957111About the history of SM can be found in [6] in the following represntations of $SU(2)_{L}$
$L_{i}$ $=$ $\left(\begin{array}[]{c}\nu_{i}\\ l_{i}\end{array}\right)_{L}\sim({\bf 1},{\bf 2},+1),\,\ l_{iR}\sim({\bf 1},{\bf 1},-2);\,$ (3)
in parenthesis it appears the transformations properties under the respective factors $(SU(3)_{C},SU(2)_{L},U(1)_{Y})$, and $i=1,2,3,$ is family index. We are omitting the quark sector and they can be found at [7, 8, 9].
In this model to generate masses for the charged leptons, we have to introduce the following scalar field $H$ in the doublet representation
$\phi$ $=$ $\left(\begin{array}[]{c}h^{+}\\ h^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1).$ (6)
This scalar fields get the following vaccum expectation value VEV
$<\phi>$ $=$ $\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v\end{array}\right),$ (9)
the scalar potential is
$V(\phi)=\mu^{2}|\phi|^{2}+\lambda|\phi|^{4},$ (10)
where $\mu$ and $\lambda$ are free parameters, when $\mu<0$ we break the $SU(2)_{L}\times U(1)_{Y}$ gauge symmetry to $U(1)_{em}$ symmetry and the minimum of the potential is
$v=\sqrt{\frac{-\mu^{2}}{\lambda}},$ (11)
when we make the following shift in the scalar field
$h^{0}=\frac{1}{\sqrt{2}}\left(v+\phi_{1}\pm\imath\phi_{2}\right),$ (12)
under this expansion, our scalar potential become
$V(\phi)=V(h^{0})-2\mu^{2}\phi^{2}_{1}+hv\phi_{1}\left(\phi^{2}_{1}+\phi^{2}_{2}\right)+\frac{h}{4}\left(\phi^{2}_{1}+\phi^{2}_{2}\right)^{2},$ (13)
where $\phi_{1}$ is CP even field and it get the following mass [^2] A nice review about it can be found in [12].
$M^{2}_{H}=-2\mu^{2},$ (14)
we must emphasize the SM do not predict any value for this mass, in fact it can get any values rom 00 until any finite values, but we can get a superior limit in its mass given by $M_{H}<1000$ GeV [10, 11]. By another hand, the field $\phi_{2}$ is CP odd field and it as massless field (Goldstone particles) and will be the responsable to give mass for $Z$ gauge boson, this mechanism is know as Brout-Englert-Higgs mechanism [12, 13]. The mass of the light Higgs boson $h^{0}$ was measured at the LHC and its value is given by [14, 15]
$M_{H}=125.09\pm 0.21\,\mathrm{(stat.)}\pm 0.11\,\mathrm{(syst.)}~{}\mathrm{GeV}.$ (15)
After this break of symmetry we get massive gauge bosons throught the terms
$\left({\cal D}\phi\right)^{\dagger}\left({\cal D}\phi\right),$ (16)
where ${\cal D}$ is the covariant derivade of the higgs fields in the SM and we get the following results [7, 8]
$M^{2}_{W}=\frac{g^{2}v^{2}}{4}=\left(80.3505\pm 0.0077\right),\,\ M^{2}_{Z}=\frac{g^{2}v^{2}}{2\cos^{2}\theta_{W}},$ (17)
where $\theta_{W}$ is the Weinberg angle and it is defined as
$e=g\sin\theta_{W}=g^{\prime}\cos\theta_{W}.$ (18)
The average values for the gauge bosnos masses are [14, 16]
$\left(M_{W}\right)_{ave}$ $=$ $\left(80.4133\pm 0.0080\right){\mbox{G}eV},$
$\left(M_{Z}\right)_{ave}$ $=$ $\left(91.1875\pm 0.0021\right){\mbox{G}eV},$ (19)
while for the Weinberg angle we have
$\sin^{2}\theta_{W}$ $=$ $0.2324\pm 0.0012$ (20)
There are also another interesting parameter to put some constrain in new physics, it is the $\rho$ parameter and it is define in the following way
$\rho\equiv\left(\frac{M^{2}_{W}}{M^{2}_{Z}\cos^{2}\theta_{W}}\right),$ (21)
in the SM its values is $1$ and it is in agreement with its experimental values given by
$\rho_{\rm EXP}=0.9998\pm 0.0008.$ (22)
The charged leptons get those masses from the following Yukawa coupling
${\cal L}^{Y}_{MP}$ $=$ $\left(g^{e}_{ij}\bar{L}_{iL}Hl_{jR}+hc\right),$ (23)
however in the SM with only left-handed neutrinos fields, $\nu_{iL}$, are massless at all perturbative order because Majorana mass term
$m_{ij}\overline{\nu^{c}_{iL}}\nu_{jL},$ (24)
break lepton number conservation and this term above can be built from the lepton doublet, defined in Eq.(3), in the following way [9]
$\overline{L^{c}_{iL}}L_{jL}\sim\left({\bf 1},{\bf 1\oplus 3},-2\right),$ (25)
and the SM respects the following accidental global symmetry [17]
$G^{global}_{SM}\equiv U(1)_{B}\times U(1)_{L_{e}}\times U(1)_{L_{\mu}}\times U(1)_{L_{\tau}},$ (26)
where $B$ is the baryon number and $L_{e},L_{\mu}$ and $L_{\tau}$ are the lepton number of each lepton flavor and it is useful strees that in the SM it does not contains, $\nu_{iR}$, sterile neutrinos [17].
There is the Fermilab’s CDF collaboration for the mass of the $W$-boson mass [18]
$\left(M_{W}\right)_{\rm CDF}=\left(80.4335\pm 0.0094\right)\,\ {\mbox{G}eV,}$ (27)
and the impact of this measurement together with Eq.(15) in the new physics beyond SM can be found in [19]. Today we know neutrinos are massive particles [20, 21, 22, 23]. One possible way for generate mass for neutrinos is to break the residual lepton number conservation.
We can break the lepton number symmetry when we introduce the new scalar $\Delta$ in the triplet representation, see Eq.(25), of the group $SU(2)$
$\Delta=\left(\begin{array}[]{cc}\Delta^{0}&\frac{h^{+}}{\sqrt{2}}\\ \frac{h^{+}}{\sqrt{2}}&H^{++}\end{array}\right)\sim\left({\bf 1},{\bf 3},2\right),\,\ \langle\Delta\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}v_{\Delta}&0\\ 0&0\end{array}\right)$ (28)
this scalar allows us to write the following Yukawa interactions [9, 24]
${\cal L}_{\Delta,L}$ $=$ $g^{\nu}_{ij}\bar{L}^{c}_{iL}\Delta L_{jL}+hc$ (29)
$=$ $g^{\nu}_{ij}\left[-\overline{\nu^{c}_{iL}}\nu_{jL}\Delta^{0}+\left(\overline{l^{c}_{iL}}\nu_{jL}-\overline{\nu^{c}_{iL}}l_{jL}\right)\frac{h^{+}}{\sqrt{2}}+\overline{l^{c}_{iL}}l_{jL}H^{++}+hc\right],$
therefore we can generate the following Majorana mass for neutrions
$M_{\rm M}\bar{\nu}^{c}_{L}\nu_{L},\,\ M_{\rm M}=g^{\nu}\frac{v_{\Delta}}{\sqrt{2}},$ (30)
it is know as Gelmini-Roncadelli scheme [24]. The mass generation mechanism of charged leptons remains the same as presented in the SM.
The masses of the gauge bosons in this case are given by [25, 26, 27]
$M^{2}_{W}$ $=$ $\frac{g^{2}}{4}\left(v^{2}+2v^{2}_{\Delta}\right),$
$M^{2}_{Z}$ $=$ $\frac{g^{2}}{4}\left(v^{2}+4v^{2}_{\Delta}\right),$ (31)
therefore we can explain the CDF new results, given in our Eq.(27), if introduce a new real triplet Higgs boson, as we shown in our Fig.(1), if $v_{\Delta}\approx 10$ GeV.
Figure 1: We show the predections for the gauge boson mass $W$ in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $v_{\Delta}$ and we show that we can explain the data from Eqs.(17,19,27).
By another hand the $\rho$ parameter in this case is given by [28]
$\rho$ $\simeq$ $1-2R\,\ R=\frac{v^{2}_{\Delta}}{v^{2}},$ (32)
and in this case the VEV must satisfy the following bound $v_{\Delta}\approx 4.92$ GeV, see our Fig.(2).
Figure 2: We show the predections for the parameter $\rho$, see our Eq.(22), in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $v_{\Delta}$.
In this scheme the scalar potential is given by [29, 30]
$V_{G}(\phi,\Delta)$ $=$ $\mu^{2}(\phi^{\dagger}\phi)+\mu^{2}_{1}Tr(\Delta^{\dagger}\Delta)+\lambda_{1}(\phi^{\dagger}\phi)^{2}+\lambda_{2}Tr(\Delta^{\dagger}\Delta)^{2}$
$+$ $\lambda_{3}(\phi^{\dagger}\phi)Tr(\Delta^{\dagger}\Delta)+\imath\lambda_{4}\epsilon_{ijk}(\phi^{\dagger}\sigma^{i}\phi)(\Delta^{j\dagger}\Delta^{k})+\lambda_{5}Tr(\Delta^{\dagger}\Delta\Delta^{\dagger}\Delta).$
Now if we make the shiftes defined in our Eq.(12), together with
$\Delta^{0}=\frac{1}{\sqrt{2}}\left(v_{\Delta}+R_{\Delta}+\imath I_{\Delta}\right).$ (34)
We get the following constraint equations
$t_{H}$ $=$ $v\left(\mu^{2}+\lambda_{1}v^{2}+\frac{\lambda_{3}}{2}v^{2}_{\Delta}-\frac{\lambda_{4}}{2}v^{2}_{\Delta}\right)=0,$
$t_{\Delta}$ $=$ $v_{\Delta}\left(\mu^{2}_{1}+\lambda_{2}v^{2}_{\Delta}+\frac{\lambda_{3}}{2}v^{2}_{\Delta}-\frac{\lambda_{4}}{2}v^{2}_{\Delta}+\lambda_{5}v^{2}_{\Delta}\right)=0.$ (35)
The mass spectrum of the neutral scalar is
$V_{quad}$ $=$ $\frac{1}{2}\left(\begin{array}[]{cc}R_{H}&R_{\Delta}\end{array}\right)M^{2}_{R}\left(\begin{array}[]{c}R_{H}\\ R_{\Delta}\end{array}\right)+\frac{1}{2}\left(\begin{array}[]{cc}I_{H}&I_{\Delta}\end{array}\right)M^{2}_{I}\left(\begin{array}[]{c}I_{H}\\ I_{\Delta}\end{array}\right).$ (42)
For the scalar of CP even, we get the following mass matrix
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda_{1}v^{2}+\frac{t_{H}}{v}&(\lambda_{3}-\lambda_{4})vv_{\Delta}\\ (\lambda_{3}-\lambda_{4})vv_{\Delta}&2\lambda_{2}v^{2}_{\Delta}+\frac{t_{\Delta}}{v_{\Delta}}\end{array}\right).$ (43)
We get two massives scalar field, the light one we will call as $h^{0}$ and the heavier as $H^{0}$. The mass of the light scalar $h$ is [^3] The mass of heavy higgs $H$ we will present below.
$m^{2}_{h}\approx 2\lambda_{1}v^{2},$ (44)
the mass $m_{h}$ as function of the VEV $v_{\Sigma}$ is shown in our Fig.(3), where we can explain the experimental. For the case of CP odd sector the mass matrix has the following representation
$M^{2}_{I}=\left(\begin{array}[]{cc}\frac{t_{H}}{v}&0\\ 0&\frac{t_{\Delta}}{v_{\Delta}}\end{array}\right).$ (45)
We get two goldstone bosons, one is eaten by $Z^{0}$ and the second is call as Majoron $M$ is $I_{\Delta}$ [24] and it is inconsistent with the experimental data [31].
Figure 3: We show the predections for the light higgs mass $h$ in the scheme of Gelmini-Roncadelli as a function of the parameter $\lambda_{1}$ and we show that we can explain the data from Eq.(15).
In this scheme the scalar potential is given by [29, 30]
$V(\phi,\Delta)$ $=$ $V_{G}(\phi,\Delta)+\mu_{M}\phi^{\dagger}\Delta\tilde{\phi},$ (46)
where $V_{G}(\phi,\Delta)$ is defined in our Eq.(LABEL:potmajorongelmini) and $\tilde{\phi}\equiv\left(\imath\sigma_{2}\phi\right)$. Under this hipotesis our Eq.(47) become
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda_{1}v^{2}&(\lambda_{3}-\lambda_{4})vv_{\Delta}+2\sqrt{2}\mu_{M}v\\ (\lambda_{3}-\lambda_{4})vv_{\Delta}+2\sqrt{2}\mu_{M}v&2\lambda_{2}v^{2}_{\Delta}-\sqrt{2}\mu_{M}\frac{v^{2}}{v_{\Delta}}\end{array}\right).$ (47)
and the mass spectrum, beyond $h$ given in Eq.(44), we get the heavy huggs $H$ and its mass is
$m^{2}_{H}\approx-\sqrt{2}\mu_{M}\frac{v^{2}}{v_{\Delta}},$ (48)
$m_{H}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(4).
Figure 4: We show the predections for the heavy Higgs mass $H$ in the scheme of Gelmini-Roncadelli, see Eq.(48), as a function of the Triplet Field VEV $V_{\Delta}$.
The mass matrix for neutral CP odd $M^{2}_{I}$ we get [29, 30]
$M^{2}_{I}=\left(\begin{array}[]{cc}-4\mu_{M}v_{\Delta}&2\mu_{M}v\\ 2\mu_{M}v&-\mu_{M}\frac{v}{v_{\Delta}}\end{array}\right).$ (49)
the majron mass
$m^{2}_{M}\approx-\sqrt{2}\mu_{M}\left(\frac{4v^{2}_{\Delta}+v^{2}}{v_{\Delta}}\right),$ (50)
$m_{M}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(5).
Figure 5: We show the predections for the Majoron mass $M$, see Eq.(50), in the scheme of Ma as a function of the Triplet Field VEV $V_{\Delta}$.
Supersymmetry, it is more known as SUSY, arose in theoretical papers more than 30 years ago independently by Golfand and Likhtman [32], Volkov and Akulov [33] and Wess and Zumino [34]. The supersymmetry algebra was introduced in [32]  [^4] They also constructed the first four-dimensional field theory with supersymmetry, (massive) quantum electrodynamics (SQED) of spinors and scalars and as suggest we will call it as Golfand-Likhtman model and they show the original version in Russian of this article received by the Editorial Board on March, 10, 1971 [35].. The Volkov-Akulov article, however, deals only with fermions and its name was “Is the Neutrino a Goldstone Particle?”. For those interested in the history of Supersymmetry see [36, 37] and for the earliest supersymmetric Standard Model see [38, 39], while anyone interested, how to work with supersymmetry see the great review article [40].
The Supersymmetric Standard Model (MSSM) is a good candidate to be the physics beyond the SM. In this model we can solve the hierarchy problem as well explaing the Higgs Masses [^5] The current status of the search for supersymmetry is presented in reference [41]. [42, 43]. In this model, as in the SM, neutrinos are massless. If we want to give masss to the neutrinos we must break $R$-Parity as well explain the new $W$ mass get from the CDF collaboration [44].
Another interesting possibility is to build Supersymmetric models with Higgs triplets as done in [45, 46, 47, 48]. There is also, an interesting proposal to explain this anomaly using the Minimal $R$-Symmetric extension of the Minimal Superymmetric Standard Model, known as MRSSM [49, 50, 51], and in this case we must respect the following inequality $|v_{\Delta}|\leq 4$ GeV [50].
Models with the following gauge symmetry
$SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ (51)
are known as $331$ for short. They are interesting possibilities for the physics at the TeV scale [52, 53, 54].
We will concentrate in the model presented in Ref. [53], and the solution to the anomaly for the mass of the $W$-boson in this model was recently presented in the following reference[55]. The scalar sector of the minimal $331$ model [53] was studied on [56, 57] the case for majoron it was studied in [25], they also show “this majoron does not mix with any other scalar fields and for this reason it does not couple, at tree level, neither to the charged leptons nor to the quarks”.
The minimal supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model, or it is more known as MSUSY331 for shortness, [58, 59, 60, 61, 62]. We have have shown the lighest higgs has mass around 125.5 GeV [60, 62]. In an recente article, we have shown we can explain the $W$-boson mass presented by the CDF in the MSUSY331 [28]. All our previous analyses in the scalar potential, it was considered only the fields $\sigma^{0}_{2}$ and $\sigma^{\prime 0}_{2}$ get VEV. Now, we want to extend those analyses presented on [25] for the MSUSY331, it meand we want to consider that both $\sigma^{0}_{1}$ and $\sigma^{\prime 0}_{1}$ get VEV in similar conditions as presented in [28].
2 Minimal Supersymmetric $331$ model.
The usual scalars fields in the minimal $331$ model are defined as $\eta,\rho,\chi$ and $S$ [53] and we can decompose those fields in $SU(2)\times U(1)$ representations in the following way [25, 63, 64, 65]:
  • ${\bf 3}_{0}\rightarrow{\bf 2}_{-1}\oplus{\bf 1}_{+2}$
    $\eta$ $=$ $\left(\begin{array}[]{c}\Phi_{\eta}\\ \eta^{+}_{2}\end{array}\right)\sim({\bf 1},{\bf 3},0),$ (54)
    $\Phi_{\eta}$ $=$ $\left(\begin{array}[]{c}\eta^{0}\\ \eta^{-}_{1}\end{array}\right)\sim({\bf 1},{\bf 2},-1),\quad\eta^{+}_{2}\sim({\bf 1},{\bf 1},+2),$ (57)
    where the singlet $\eta^{+}_{2}$ was proposed by Zee [66]
  • ${\bf 3}_{1}\rightarrow{\bf 2}_{1}\oplus{\bf 1}_{4}$
    $\rho$ $=$ $\left(\begin{array}[]{c}\Phi_{\rho}\\ \rho^{++}\end{array}\right)\sim({\bf 1},{\bf 3},+1),$ (60)
    $\Phi_{\rho}$ $=$ $\left(\begin{array}[]{c}\rho^{+}\\ \rho^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1),\quad\rho^{++}\sim({\bf 1},{\bf 1},+4)$ (63)
  • ${\bf 3}_{-1}\rightarrow{\bf 2}_{-3}\oplus{\bf 1}_{0}$
    $\chi$ $=$ $\left(\begin{array}[]{c}\Phi_{\chi}\\ \chi^{0}\end{array}\right)\sim({\bf 1},{\bf 3},-1),$ (66)
    $\Phi_{\chi}$ $=$ $\left(\begin{array}[]{c}\chi^{-}\\ \chi^{--}\end{array}\right)\sim({\bf 1},{\bf 2},-3),\quad\chi^{0}\sim({\bf 1},{\bf 1},0)$ (69)
  • ${\bf 6}_{0}\rightarrow{\bf 3}_{2}\oplus{\bf 2}_{-1}\oplus{\bf 1}_{-4}$
    $S$ $=$ $\left(\begin{array}[]{cc}T&\frac{\Phi_{S}}{\sqrt{2}}\\ \frac{\Phi^{T}_{S}}{\sqrt{2}}&H^{--}_{2}\end{array}\right)\sim({\bf 1},{\bf 6}^{*},0),$ (72)
    $\Phi_{S}$ $=$ $\left(\begin{array}[]{c}h^{+}_{2}\\ \sigma^{0}_{2}\end{array}\right)\sim({\bf 1},{\bf 2},-1),\quad H^{--}_{2}\sim({\bf 1},{\bf 1},-4),$ (75)
where $\Phi_{S}$ is the Higgs doublet boson of SM, see Eq.(12), and the following triplet
$T$ $=$ $\left(\begin{array}[]{cc}\sigma^{0}_{1}&\frac{h^{+}_{1}}{\sqrt{2}}\\ \frac{h^{+}_{1}}{\sqrt{2}}&H^{--}_{1}\\ \end{array}\right)\sim({\bf 1},{\bf 3},+2),$ (78)
where $T$ is the Gelmini-Rocandelli $SU(2)$ triplet, see Eq.(28), and $H^{--}_{2}$ appear in the model of Babu presented in reference [67]. We redefine the neutral components of the scalars defined above in the following way
$\eta^{0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\eta}+R^{0}_{\eta}+\imath I^{0}_{\eta}\right),\quad\rho^{0}=\frac{1}{\sqrt{2}}\left(v_{\rho}+R^{0}_{\rho}+\imath I^{0}_{\rho}\right),$
$\chi^{0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\chi}+R^{0}_{\chi}+\imath I^{0}_{\chi}\right),\quad\sigma^{0}_{1}=\frac{1}{\sqrt{2}}\left(v_{\sigma^{0}_{1}}+R^{0}_{\sigma^{0}_{1}}+\imath I^{0}_{\sigma^{0}_{1}}\right),$
$\sigma^{0}_{2}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\sigma^{0}_{2}}+R^{0}_{\sigma^{0}_{2}}+\imath I^{0}_{\sigma^{0}_{2}}\right).$ (79)
In order to implement supersymmetry and also at same time to cancel chiral anomalies we must to introduce the following scalars fields [60, 61]
$\eta^{\prime}$ $=$ $\left(\begin{array}[]{c}\eta^{\prime 0}\\ \eta^{\prime+}_{1}\\ \eta^{\prime-}_{2}\end{array}\right)\sim({\bf 1},{\bf 3}^{*},0),\quad\rho^{\prime}=\left(\begin{array}[]{c}\rho^{\prime-}\\ \rho^{\prime 0}\\ \rho^{\prime--}\end{array}\right)\sim({\bf 1},{\bf 3}^{*},-1),$ (86)
$\chi^{\prime}$ $=$ $\left(\begin{array}[]{c}\chi^{\prime+}\\ \chi^{\prime++}\\ \chi^{\prime 0}\end{array}\right)\sim({\bf 1},{\bf 3}^{*},+1),\quad S^{\prime}=\left(\begin{array}[]{ccc}\sigma^{\prime 0}_{1}&\frac{h^{\prime-}_{2}}{\sqrt{2}}&\frac{h^{\prime+}_{1}}{\sqrt{2}}\\ \frac{h^{\prime-}_{2}}{\sqrt{2}}&H^{\prime--}_{1}&\frac{\sigma^{\prime 0}_{2}}{\sqrt{2}}\\ \frac{h^{\prime+}_{1}}{\sqrt{2}}&\frac{\sigma^{\prime 0}_{2}}{\sqrt{2}}&H^{\prime++}_{2}\end{array}\right)\sim({\bf 1},{\bf 6},0).$ (93)
again we redefine
$\eta^{\prime 0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\eta^{\prime}}+R^{0}_{\eta^{\prime}}+\imath I^{0}_{\eta^{\prime}}\right),\quad\rho^{\prime 0}=\frac{1}{\sqrt{2}}\left(v_{\rho^{\prime}}+R^{0}_{\rho^{\prime}}+\imath I^{0}_{\rho^{\prime}}\right),$
$\chi^{\prime 0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\chi^{\prime}}+R^{0}_{\chi^{\prime}}+\imath I^{0}_{\chi^{\prime}}\right),\quad\sigma^{\prime 0}_{1}=\frac{1}{\sqrt{2}}\left(v_{\sigma^{\prime 0}_{1}}+R^{0}_{\sigma^{\prime 0}_{1}}+\imath I^{0}_{\sigma^{\prime 0}_{1}}\right),$
$\sigma^{\prime 0}_{2}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\sigma^{\prime 0}_{2}}+R^{0}_{\sigma^{\prime 0}_{2}}+\imath I^{0}_{\sigma^{\prime 0}_{2}}\right).$ (95)
Our superpotencial is defined as [60, 61, 62]
$W=\frac{W_{2}}{2}+\frac{W_{3}}{3}+hc,$ (96)
where
$W_{2}$ $=$ $\sum_{i=1}^{3}\mu_{0i}(\hat{L}_{i}\hat{\eta}^{\prime})+\mu_{\eta}(\hat{\eta}\hat{\eta}^{\prime})+\mu_{\rho}(\hat{\rho}\hat{\rho}^{\prime})+\mu_{\chi}(\hat{\chi}\hat{\chi}^{\prime})+\mu_{S}Tr[(\hat{S}\hat{S}^{\prime})],$
$W_{3}$ $=$ $\sum_{i=1}^{3}\left[\sum_{j=1}^{3}\left(\sum_{k=1}^{3}\lambda_{1ijk}(\epsilon\hat{L}_{i}\hat{L}_{j}\hat{L}_{k})+\lambda_{2ij}(\epsilon\hat{L}_{i}\hat{L}_{j}\hat{\eta})+\lambda_{3ij}(\hat{L}_{i}\hat{S}\hat{L}_{j})\right)+\lambda_{4i}(\epsilon\hat{L}_{i}\hat{\chi}\hat{\rho})\right]$ (97)
$+$ $f_{1}(\epsilon\hat{\rho}\hat{\chi}\hat{\eta})+f_{2}(\hat{\chi}\hat{S}\hat{\rho})+f_{3}(\hat{\eta}\hat{S}\hat{\eta})+f_{4}\epsilon_{ijk}\epsilon_{lmn}\hat{S}_{il}\hat{S}_{jm}\hat{S}_{kn}+f^{\prime}_{1}(\epsilon\hat{\rho}^{\prime}\hat{\chi}^{\prime}\hat{\eta}^{\prime})+f^{\prime}_{2}(\hat{\chi}^{\prime}\hat{S}^{\prime}\hat{\rho}^{\prime})$
$+$ $f^{\prime}_{3}(\hat{\eta}^{\prime}\hat{S}^{\prime}\hat{\eta}^{\prime})+f^{\prime}_{4}\epsilon_{ijk}\epsilon_{lmn}\hat{S}^{\prime}_{il}\hat{S}^{\prime}_{jm}\hat{S}^{\prime}_{kn}+\sum_{i=1}^{3}\left[\kappa_{1i}(\hat{Q}_{3}\hat{\eta}^{\prime})\hat{u}^{c}_{i}+\kappa_{2i}(\hat{Q}_{3}\hat{\rho}^{\prime})\hat{d}^{c}_{i}\right]+\kappa_{3}(\hat{Q}_{3}\hat{\chi}^{\prime})\hat{J}^{c}$
$+$ $\sum_{\alpha=1}^{2}\left[\sum_{i=1}^{3}\left(\kappa_{4\alpha i}(\hat{Q}_{\alpha}\hat{\eta})\hat{d}^{c}_{i}+\kappa_{5\alpha i}(\hat{Q}_{\alpha}\hat{\rho})\hat{u}^{c}_{i}\right)+\sum_{\beta=1}^{2}\kappa_{6\alpha\beta}(\hat{Q}_{\alpha}\hat{\chi})\hat{j}^{c}_{\beta}+\sum_{i=1}^{3}\sum_{j=1}^{3}\kappa_{7\alpha ij}(\hat{Q}_{\alpha}\hat{L}_{i})\hat{d}^{c}_{j}\right]$
$+$ $\sum_{i=1}^{3}\left\{\sum_{j=1}^{3}\left[\sum_{k=1}^{3}\xi_{1ijk}\hat{d}^{c}_{i}\hat{d}^{c}_{j}\hat{u}^{c}_{k}+\sum_{\beta=1}^{2}\left(\xi_{2ij\beta}\hat{u}^{c}_{i}\hat{u}^{c}_{j}\hat{j}^{c}_{\beta}\right]+\xi_{3i\beta}\hat{d}^{c}_{i}\hat{J}^{c}\hat{j}^{c}_{\beta}\right)\right\}.$
The terms proportional for $f_{2},f_{4},f^{\prime}_{2}$ and $f^{\prime}_{4}$ violate the conservation of the
${\cal F}=B+L,$ (98)
quantum number [60, 61, 62].
As we said, we have shown we can explain the $W$-boson mass presented by the CDF in the MSUSY331 in two possible ways [28]. For somephenomenological analyses is useful define the following parameter [68, 69] [^6] We want to perform similar constraints as presented in [68, 69] in this model.
$\xi_{WW^{\prime}}=\left(\frac{M_{W}}{M_{W^{\prime}}}\right)^{2},\,\ \xi_{ZZ^{\prime}}=\left(\frac{M_{Z}}{M_{Z^{\prime}}}\right)^{2}$ (99)
  • 1-)
    $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV and $v_{\sigma_{2}}=10\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=11.19\,\ {\rm GeV}$ and the parameters defined in our Eq.(99) is shown in our Tab.(1).
    $v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
    1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
    1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
    2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
    2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
    3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
    3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
    4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
    4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
    5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
    Table 1: The parameters $\xi_{WW^{\prime}}$ and $\xi_{ZZ^{\prime}}$, see Eq.(99), as function of $v_{\chi}$ in GeV and we take $v_{\chi^{\prime}}=1000$ GeV and the others vev are given in Eqs.(63,64) and Eq.(73) [28].
  • 2-)
    Several solution as we shown in our Tab.(2). We will use the following point for our numerical analyses $v_{\sigma_{1}}=4.89\,\ {\rm GeV},v_{\sigma_{2}}=10\,\ {\rm GeV},v_{\sigma^{\prime}_{1}}=0.5\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=8.77\,\ {\rm GeV}$ and the parameters defined in our Eq.(99) is shown in our Tab.(3).
$v_{\sigma_{1}}$ GeV $v_{\sigma^{\prime}_{1}}$ GeV
0.5 4.89453
1.0 4.81730
1.5 4.68577
2.0 4.49515
2.5 4.2375
3.0 3.89954
3.5 3.4578
4.0 2.86468
4.5 1.98907
Table 2: Defined some values of $v_{\sigma_{1}}$ we show the numerical values for $v_{\sigma^{\prime}_{1}}$ GeV take of our study presented in [28].
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
Table 3: The parameters $\xi_{WW^{\prime}}$ and $\xi_{ZZ^{\prime}}$, see Eq.(99), as function of $v_{\chi}$ in GeV and we take $v_{\chi^{\prime}}=1000$ GeV and the others vev are given in Eqs.(63,64) and Eq.(73) [28].
3 Case 1
In this case, as we choose $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV, we got the $8\times 8$ mass matrix as presented in one preliminar numerical analyses, presented in [62]. We will use below the following set of parameters in the scalar potential:
$f_{1}=1.12,\quad f_{3}=1.03,\quad f^{\prime}_{1}=f^{\prime}_{3}=10^{-6},\quad{\rm(dimensionless)}$ (100)
and
[…]
we also use the constraint
$V^{2}_{\eta}+V^{2}_{\rho}+2V^{2}_{2}=(246\;{\rm GeV})^{2}$ (102)
coming from $M_{W}$, where, we have defined $V^{2}_{\eta}=v^{2}_{\eta}+v^{\prime 2}_{\eta}$ and $V^{2}_{\rho}=v^{2}_{\rho}+v^{\prime 2}_{\rho}$ and $V^{2}_{2}=v^{2}_{\sigma_{2}}+v^{2}_{\sigma^{\prime}_{2}}$. Assuming that $v_{\eta}=20$, $v_{\chi}=1500$, $v_{\sigma_{2}}=10$, $v^{\prime}_{\eta}=v^{\prime}_{\rho}=v^{\prime}_{\sigma_{2}}=v^{\prime}_{\chi}=1$ in GeV, the value of $v_{\rho}$ is fixed by the constraint above. We get similar values are presented in [62]. We present in our Fig.(6) our results for the masses of $M_{h}$ and $M_{A}$ and we see it is in agreement with Eq.(15).
Figure 6: The masses of light scalars, $M_{h}$ and light pseudoscalar $M_{A}$, in terms of $v_{\chi}$ for $v_{\chi^{\prime}}=10$ GeV and $M^{exp}_{h}$ is the experimental values given at our Eq.(15).
4 Case 2
Here we got two $10\times 10$ mass matrix. The bases for our scalars of CP even is
$\left(H_{\eta},H_{\rho},H_{\chi},H_{\sigma_{1}},H_{\sigma_{2}},H_{\eta^{\prime}},H_{\rho^{\prime}},H_{\chi^{\prime}},H_{\sigma^{\prime}_{1}},H_{\sigma^{\prime}_{2}}\right)^{T},$ (103)
while the base for our scalars of CP odd is
$\left(F_{\eta},F_{\rho},F_{\chi},F_{\sigma_{1}},F_{\sigma_{2}},F_{\eta^{\prime}},F_{\rho^{\prime}},F_{\chi^{\prime}},F_{\sigma^{\prime}_{2}}\right)^{T}.$ (104)
Our results are shown in Figs.(7,8,9,10) where we can say
  • $M_{h}\sim 125.5\,\ {\mbox{G}eV}$;
  • $M_{A}>150\,\ {\mbox{G}eV}$,
and here we used $f_{2}=f_{4}=f^{\prime}_{2}=f^{\prime}_{4}=0$ and in the sector of CP odd we got three goldstone bosons. One of them is the Majoron presented at [25].
Figure 7: The masses of light scalars, $M_{h}$, in terms of $v_{\chi}$ for for some values of $v_{\chi^{\prime}}$ shown at box GeV and $M^{exp}_{h}$ is the experimental values given at our Eq.(15).
Figure 8: The masses of light scalars, $M_{A}$, in terms of $v_{\chi}$ for for some values of $v_{\chi^{\prime}}$ shown at box.
Figure 9: The masses of light scalars, $M_{h}$, in terms of $v_{\chi^{\prime}}$ for for some values of $v_{\chi}$ shown at box GeV and $M^{exp}_{h}$ is the experimental values given at our Eq.(15).
Figure 10: The masses of light scalars, $M_{A}$, in terms of $v_{\chi^{\prime}}$ for for some values of $v_{\chi}$ shown at box.
However if we take $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ not no null the Majoron dissapear but it has mass of few GeV, the same results as presented in [25], but the values of $M_{A}$ is near the ones we present in our Figs.(7,9).
5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet not no vanishing and we shown our light scalar fields has mass near the value $M_{h}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(15), and $M_{A}>150\,\ {\mbox{G}eV}$.
Acknowledgments
We would like thanks V. Pleitez for useful discussions above $331$ models. We also to thanks IFT for the nice hospitality during my several visit to perform my studies about the severals $331$ Models and also for done this article.